OpenStudy (anonymous):

find the area of the trapezoid h=5 base=10 w=4

Parth (parthkohli):

\[\rm Area~of~a~trapezoid = \dfrac{h}{2}\left(s_1 + s_2\right)\]

OpenStudy (anonymous):

i don't know how to do that

Parth (parthkohli):

Is \(w\) a side here?

OpenStudy (anonymous):

w is the length inside of the trapezoid i think its the width

Parth (parthkohli):

Oh, the width. Then it should be \(h\times w\).

OpenStudy (anonymous):

so when i multiply the w and h ill get the answer

Parth (parthkohli):

Yeah.

OpenStudy (anonymous):

so the answer is 20